a 2ft times 5ft times 2in plate of steel will weight......(answers are 816lbs,408lbs,and 204lbs) the formulsa giver are for Volume=HxWxL and weight=volume x unit weight of material. the weight of steel is 490lbs/cubic ft. i don't want you just to give me the answer because then I haven't learned how to do it myself, can you pleaseeee show me step by step what to do I am so lost here
1
Expert's answer
2011-10-07T08:18:15-0400
We have that 1 foot = 12 inches, so the volume of steel in cubic foots
is

& 2ft x& 5ft x 2in = 2 x 5 x 2/12 = 20/12 = 1.66667 cubic
foots.

Therefore the weigth of that plate is

& W = 490 *
1.666667 = 816,6666 lbs